### v3

Abstract
Many practical optimization problems lack strong convexity. Fortunately, recent studies have revealed that first-order algorithms also enjoy linear convergences under various weaker regularity conditions. While the relationship among different conditions for convex and smooth functions is well-understood, it is not the case for the nonsmooth setting. In this paper, we go beyond convexity and smoothness, and clarify the connections among common regularity conditions in the class of weakly convex functions, including strong convexity, restricted secant inequality, subdifferential error bound, Polyak-Łojasiewicz inequality, and quadratic growth. In addition, using these regularity conditions, we present a simple and modular proof for the linear convergence of proximal point method (PPM) for convex and weakly convex optimization problems. The linear convergence also holds when the subproblems of PPM are solved inexactly with a proper control of inexactness.
keywords:
Error bound, Polyak-Łojasiewicz inequality, Quadratic growth, Proximal point method, Linear convergence, Weakly convex functions
1 Introduction
With the success of machine learning, (sub)gradient-type algorithms, such as (sub)gradient descent, coordinate descent, Frank-Wolf algorithm, and other variants, have gained much attention over the past decades [beck2017first, bottou2018optimization]. For smooth and/or convex cases, (sub)gradient methods are most well-understood. It is well-known that the basic gradient descent algorithm achieves linear convergence for minimizing smooth and strongly convex functions [nesterov2018lectures]. However, strong convexity is often not satisfied in practice, and many fundamental models in machine learning lack this good property [agarwal2010fast].
This motivates the study of different regularity conditions that are weaker than strong convexity, such as Polyak-Łojasiewicz inequality [polyak1963gradient], restricted secant inequality [zhang2013gradient], subdifferential error bound [drusvyatskiy2018error], and quadratic growth [necoara2019linear]. These regularity conditions go beyond strong convexity, and they mainly involve the relationship concerning the distance to the solution set, the cost value gap, and the length of the (sub)gradient. They often play a crucial step in many algorithm analyses. It has been shown that under those regularity conditions, a range of first-order optimization algorithms still enjoy linear convergence in many cases; see e.g., [polyak1963gradient, karimi2016linear, cui2019r, drusvyatskiy2018error, diaz2023optimal, garber2023linear, liao2023overview, ding2023revisiting]. Moreover, these convergence analyses even become more elegant under suitable regularity conditions; examples include the gradient descent method for smooth optimization [nesterov2018lectures, Theorem 2.1.15]; the proximal gradient method for nonsmooth convex optimization [drusvyatskiy2018error, Theorem 3.2]; and the proximal point method (PPM) to solve $0\in T(x)$ where $T$ is a maximal monotone set value mapping [leventhal2009metric].
Recent studies have revealed the relationship between some regularity conditions under various assumptions on the smoothness and convexity of the objective function [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth, guille2021study, rebjock2023fast]. We postpone the comparison between different studies after we formally define regularity conditions (see Remark 1). Different conditions are useful to prove the convergence of various algorithms. For example, the subdifferential error bound yields a clean proof for the proximal point method (PPM) [rockafellar1976monotone, leventhal2009metric]; the step-length type error bound becomes useful for the proximal gradient method [drusvyatskiy2018error]; Polyak-Łojasiewicz inequality leads to a clean proof of the gradient descent [polyak1963gradient, karimi2016linear]; quadratic growth plays an important role in the analysis of proximal bundle methods [diaz2023optimal] and spectral bundle methods [ding2023revisiting, liao2023overview]. In some cases, it may be easier to establish one regularity condition than another one and then use the equivalence between the two to derive convergence results, e.g. [cui2016asymptotic, Proposition 3.3]. Therefore, understanding the relationship or equivalence between common regularity conditions becomes crucial in algorithm analyses.
Despite the results mentioned above on the equivalence [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth, guille2021study, rebjock2023fast], their scope remains limited to one of the following scenarios: 1) smooth functions, or 2) nonsmooth but convex functions. While smooth and convex (but not strongly convex) problems cover a variety of applications, practical applications routinely deal with problems that lack both qualities (e.g., training nonsmooth and nonconvex deep networks). Recent studies have further identified another amenable problem class: weakly convex functions [davis2019stochastic]. This class of functions includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning [drusvyatskiy2020subgradient, atenas2023unified]. Some algorithms for convex functions can also be extended to solve weakly convex functions, e.g., the classical bundle method [liang2023proximal], the stochastic projected subgradient method [davis2019proximally], and the damped proximal augmented Lagrangian method [dahal2023damped]. A recent study also proposes a unified analysis framework to analyze the convergence of algorithms designed for weakly convex functions [atenas2023unified]. In this paper, we aim to generalize the previous results on the equivalence of different regularity conditions [karimi2016linear, drusvyatskiy2018error, artacho2008characterization, bolte2017error, ye2021variational, zhu2023unified, drusvyatskiy2021nonsmooth] from the class of smooth functions (and the class of convex functions) to the class of nonsmooth weakly convex functions. These findings will benefit algorithm analyses from various perspectives.
To demonstrate the equivalence between different regularity conditions, we revisit the classical (inexact) PPM, a conceptually simple algorithm that serves as a guideline for many algorithm developments [drusvyatskiy2017proximal]. In particular, we provide a simple new proof for the linear convergence in both the convex and weakly convex settings using Polyak-Łojasiewicz inequality, subdifferential error bound, and quadratic growth. Our proof is partially inspired by a recent result on spectral bundle methods [ding2023revisiting, liao2023overview] and the analysis in [leventhal2009metric]. As a by-product of our investigation, we clarify some subtleties in inexact PPM in terms of the inexactness and feasibility, which are less emphasized in the literature. In this paper, our contributions are as follows.
  • We first clarify the relationship among common regularity conditions, including strongly convex, restricted secant inequality, subdifferential error bound, Polyak-Łojasiewicz inequality, and quadratic growth, in the class of weakly convex functions (Theorem 3.1). Most of the proofs follow from some elementary analysis. The connection between Polyak-Łojasiewicz inequality and subdifferential error bound utilizes the notion of slope in [drusvyatskiy2021nonsmooth, Section 2] and requires some sophisticated arguments.We also derive the coefficient conversion among different conditions (see Table 1).
  • Building upon the equivalence for different regularity conditions, we present a simple and modular proof for linear convergence of the PPM for convex optimization [rockafellar1976monotone] (Theorem 4.2). The linear convergence also extends to the class of weakly convex functions with a proper initialization (Theorem 4.3). Our linear convergence result allows non-unique minimizers which is more general than a classical assumption of local Lipschitz continuity in [rockafellar1976monotone] and also relaxes an upper Lipschitz condition in [luque1984asymptotic].
  • Finally, we clarify some subtleties in the inexact PPM regarding inexactness stopping criteria and feasibility (Section 5.1). We further establish the linear convergence of inexact PPM (Theorem 5.3) under standard assumptions on the inexactness of the subproblem. Our proof is clean and modular, which combines a key inequality in [luque1984asymptotic] with Theorem 4.2.
A preliminary summary of this work appeared in [pmlr-v242-liao24a]. In this paper, we 1) complete the detailed proofs for all theorems and derive coefficient conversion among different conditions in Table 1, 2) extend the linear convergence result of the PPM from optimization on convex functions to weakly convex functions (Theorem 4.3), and 3) provide extra technical discussions on inexact PPM (Section 5.1). The rest of this paper is structured as follows. Section 2 summarizes the linear convergence of gradient descent as motivation and introduces some preliminaries in nonsmooth optimization techniques. Section 3 presents the relationship among different regularity conditions. Sections 4 and 5 focus on the (inexact) PPM and establish the sublinear and linear convergences. Section 6 presents three numerical experiments. Section 7 concludes this paper. Some auxiliary proofs and discussions are provided in Appendices D, A, D, C and B.
[…]
2 Motivation and preliminaries
In this section, we first motivate the study of equivalent regularity conditions by revisiting the standard gradient descent algorithm, and then review some techniques in nonsmooth optimization.
2.1 Motivation: Linear convergence of gradient descent algorithms
To motivate our discussion, consider a smooth convex optimization problem $\min_{x}\,f(x)$, where $f:\mathbb{R}^{n}\to{\mathbb{R}}$ is a convex and $L$-smooth function, i.e., its gradient is $L$-Lipschitz satisfying $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|,\;\forall x,$ $y\in\mathbb{R}^{n}.$ Let $S:=\operatorname*{\arg\!\min}_{x}\,f(x)$ be the set of optimal solutions. Assume $S\neq\emptyset$ and denote $f^{\star}=\min_{x}f(x)$. The standard gradient descent (GD) follows the update
$x_{k+1}=x_{k}-t_{k}\nabla f(x_{k}),$ (1)
where $t_{k}>0$ is the step size. A textbook result says that when choosing a constant step size $t_{k}=\frac{1}{L}$, the GD iterates converge to $f^{\star}$ with a sublinear rate (precisely, $f(x_{k})-f^{\star}\leq L\mathrm{dist}^{2}(x_{0},S)/(2k)$); see e.g., [nesterov2018lectures, Corollary 2.1.2]. If the function $f$ is strongly convex, GD achieves a global linear convergence [nesterov2018lectures, Theorem 2.1.15].
However, the assumption of strong convexity is often not satisfied in practice. It is known that some alternative weaker assumptions are sufficient for linear convergences. We here introduce two notions: restricted secant inequality (RSI) [zhang2013gradient], and Polyak-Łojasiewicz (PL) inequality [polyak1963gradient]. A differentiable function $f$ satisfies RSI if there exists $\mu_{\mathrm{r}}>0$ such that
$\mathopen{}\mathclose{{\left\langle\nabla f(x),x-{x}^{\star}}}\right\rangle\geq\mu_{\mathrm{r}}\|x-{x}^{\star}\|^{2}=\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S),\forall x\in\mathbb{R}^{n},{x}^{\star}\in\Pi_{S}(x),$ (2)
and it satisfies the PL inequality if there exists $\mu_{\mathrm{p}}>0$ such that
$\|\nabla f(x)\|^{2}\geq 2\mu_{\mathrm{p}}(f(x)-f^{\star}),\;\forall x\in\mathbb{R}^{n}.$ (3)
Note that both RSI 2 and PL 3 imply that any stationary point of $f$ is a global minimum. However, they do not imply the uniqueness of stationary points or the convexity of the function. One can think that RSI 2 (resp. PL 3) requires that the gradient $\nabla f(x)$ grows faster than that of quadratic functions when moving away from the solution set $S$ (resp. the optimal value $f^{\star}$). Linear convergence of GD under the PL inequality was first proved in [polyak1963gradient], and linear convergence under RSI was discussed in [guille2022gradient, Proposition 1] and [zhang2020new, Proposition 1]. We summarize a simple version below.
Theorem 2.1 (Linear convergence of GD).
Consider the optimization problem $\min_{x\in\mathbb{R}^{n}}f(x),$ where $f$ is an $L$-smooth function. Suppose its solution set $S$ is nonempty. If RSI 2 holds with $\mu_{\mathrm{r}}>0$ and PL inequality 3 holds with $\mu_{\mathrm{p}}>0$, then the GD algorithm 1 with a constant stepsize $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$ has global linear convergence for iterates and function values, i.e.,
$\mathrm{dist}(x_{k+1},S)$ $\leq\omega_{1}\cdot\mathrm{dist}(x_{k},S),$ (4a)
$f(x_{k+1})-f^{\star}$ $\leq\omega_{2}\cdot(f(x_{k})-f^{\star}),$ (4b)
with $\omega_{1}=\sqrt{1-\mu_{\mathrm{r}}^{2}/L^{2}}\in(0,1)$ and $\omega_{2}=(L^{3}-2\mu_{\mathrm{r}}L\mu_{\mathrm{p}}+\mu_{\mathrm{r}}^{2}\mu_{\mathrm{p}})/L^{3}\in(0,1)$.
Thanks to RSI 2 and PL inequality 3, the proof of Theorem 2.1 is very elegant, taking only a few lines. For self-completeness, we provide a simple proof and some additional discussions in Section B.1. In particular, the RSI 2 leads to a quick proof of 4a, and the PL 3 allows for a simple proof of 4b. It is known that for $L$-smooth convex functions, the two conditions RSI 2 and PL 3 are equivalent (cf. [karimi2016linear, Theorem 2]). Some recent studies, such as [bolte2017error, necoara2019linear, zhang2017restricted], have explored the relationships among different regularity conditions for linear convergences. A nice summary appeared in [karimi2016linear, Theorem 2], but it only works in the context of $L$-smooth functions.
In this paper, we aim to characterize the relationships among different regularity conditions for nonsmooth and nonconvex functions (Section 3), and apply them to derive simple and clean proofs for linear convergences of (inexact) proximal point methods for convex and weakly convex (possibly nonsmooth) optimization (Sections 4 and 5). Before that, we next introduce some key techniques for analyzing nonsmooth problems.
2.2 The notion of slope and Ekeland’s variational principle
In this subsection, we review three important techniques in nonsmooth optimization, 1) the notion of slope, 2) a key distance estimation result, and 3) Ekeland’s variational principle, which will play an important role in our proofs later.
We first introduce the Fréchet subdifferential. Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed function. We define the Fréchet subdifferential (see e.g., [li2020understanding, Page 27]):
$\hat{\partial}f(x):=\mathopen{}\mathclose{{\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle}{\|y-x\|}\geq 0}}\right\}.$
If $f$ is convex, the Fréchet subdifferential $\hat{\partial}f$ is the same as the usual convex subdifferential, i.e.,
$\hat{\partial}f(x)=\partial f(x):=\{s\in\mathbb{R}^{n}\mid f(y)\geq f(x)+\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle,\forall y\in\mathbb{R}^{n}\},\;\;\forall x\in\mathbb{R}^{n},$
and if $f$ is differentiable, Fréchet subdifferential reduces to the usual gradient, i.e., $\hat{\partial}f(x)=\{\nabla f(x)\},\forall x\in\mathbb{R}^{n}$.
Definition 1 (Slope [drusvyatskiy2021nonsmooth, Section 2]).
Consider a closed function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $\bar{x}$ with $f(\bar{x})$ finite. The slope of $f$ at $\bar{x}$ is its maximal instantaneous rate of decrease:
$|\nabla f|(\bar{x}):=\limsup_{x\to\bar{x}}\frac{(f(\bar{x})-f(x))^{+}}{\|x-\bar{x}\|},$
where we use the notation $r^{+}=\max\{0,r\}$.
If $f$ is smooth, the slope $|\nabla f|(\bar{x})$ is simply the norm of its gradient $\|\nabla f(\bar{x})\|$. If $f(x)$ is convex, the slope $|\nabla f|(\bar{x})$ is equivalent to the length of the minimal norm element in the subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x}))$. If it is $\rho$-weakly convex (i.e., $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex), the slope $|\nabla f|(\bar{x})$ is the same as the length of the minimal norm element in the Fréchet subdifferential, i.e., $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$ (see [ioffe2017variational, drusvyatskiy2013slope] for more details). We summarize these useful properties into a lemma below; a brief proof is presented in Section B.2 for self-completeness:
Lemma 2.1.
Consider a closed function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $\bar{x}\in\mathbb{R}^{n}$ with $f(\bar{x})$ finite. The following statements for the slope hold.
  1. 1.
    If $f$ is smooth, then $|\nabla f|(\bar{x})=\|\nabla f(\bar{x})\|$;
  2. 2.
    If $f$ is convex, then $|\nabla f|(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x}))$;
  3. 3.
    If $f$ is $\rho$-weakly convex, then $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$.
We then introduce a key result from [drusvyatskiy2015curves, Lemma 2.5] that will serve as a crucial step to estimate the distance to the optimal solution set.
Lemma 2.2 (Lemma 2.5 [drusvyatskiy2015curves]).
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed function. Suppose for some point $x\in\text{dom}(f)$, there are constants $\alpha<f(x)$ and $r>0,K>0$ such that $f(x)-\alpha<Kr$ and
$|\nabla f|(u)\geq r,\qquad\forall u\in[\alpha<f\leq f(x)],\|u-x\|\leq K.$
Then the sublevel set $f\leq\alpha]$ is nonempty and $\mathrm{dist}(x,[f\leq\alpha])\leq(f(x)-\alpha)/r.$
Note that if $\alpha=f^{\star}=\inf_{x\in\mathbb{R}^{n}}f(x)$ in Lemma 2.2, we can upper bound the distance to the solution set $\mathrm{dist}(x,[f=f^{\star}])$ by $(f(x)-f^{\star})/r$.
Finally, we review Ekeland’s variational principle [ekeland1974variational].
Theorem 2.2 (Ekeland’s variational principle [ekeland1974variational]).
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed function. Suppose $\epsilon>0$ and $z\in[f\leq\inf_{x\in\mathbb{R}^{n}}f(x)+\epsilon]$. Then for any $\rho>0$ there exists $y\in\mathbb{R}^{n}$ such that
$\|z-y\|$ $\leq\epsilon/\rho,$ (5a)
$f(y)$ $\leq f(z),$ (5b)
$f(x)+\rho\|x-y\|$ $>f(y),\;\;\forall x\in\mathbb{R}^{n}/\{y\}.$ (5c)
Note that 5a means that the distance between $y$ and $z$ is bounded, 5b tells us that $y$ is also in the sublevel set, and 5c shows that $y$ uniquely minimizes the function $f+\rho\|\cdot-y\|$.
From the definition of slope (Definition 1), one immediate consequence of Theorem 2.2 is that the slope of $y$ is also bounded by $\rho$ as
$|\nabla f|(y)=\limsup_{x\to y}\frac{(f(y)-f(x))^{+}}{\|x-y\|}\leq\limsup_{x\to y}\frac{\rho\cdot\mathrm{dist}(x,y)}{\|x-y\|}=\rho.$ (6)
Lemma 2.2 and 2.2 will play an important role in our analysis of different regularity conditions in Theorem 3.1.
3 Relationships between regularity conditions
In this section, we aim to show the equivalence between different regularity conditions in the class of weakly convex (potentially nonsmooth) functions. A function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is called $\rho$-weakly convex if the function $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex. The class of weakly convex functions is very rich: it includes all convex functions, $L$-smooth functions, certain compositions of convex functions with smooth functions, and many cost functions in modern machine learning applications; we refer interested readers to [drusvyatskiy2020subgradient, atenas2023unified] for more details.
3.1 Regularity conditions and their relationship
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper, closed, $\rho$-weakly convex function. Let $S$ be the optimal solution set of $f$, i.e., $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$, and we assume $S\neq\emptyset$. Let $f^{\star}=\min_{x\in\mathbb{R}^{n}}f(x)$ and $\nu>0$. We consider the following five regularity conditions:
  1. 1.
    Local Strong Convexity (SC): there exists a positive constant $\mu_{\mathrm{s}}>0$ such that
    $f(x)+\langle g,y-x\rangle+\frac{\mu_{\mathrm{s}}}{2}\|y-x\|^{2}$ $\leq f(y),\quad\forall x,y\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x).$ (SC)
  2. 2.
    Restricted Secant Inequality (RSI): there exists a positive constant $\mu_{\mathrm{r}}>0$ such that
    $\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S)$ $\leq\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle,\quad\forall x\in[f\leq f^{\star}+\nu],g\in\hat{\partial}f(x),\hat{x}\in\Pi_{S}(x).$ (RSI)
  3. 3.
    Error bound (EB) [^2] Error bound is closely related to the notion of metric subregularity at an optimal point $x^{\star}$ for $0\in\hat{\partial}f(x^{\star})$ [artacho2008characterization, Definition 2.3]: there exist a constant $\alpha>0$ and a set $\mathcal{U}$ containing $x^{\star}$ such that $\mathrm{dist}(x,(\hat{\partial}f)^{-1}(0))\leq\alpha\cdot\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in\mathcal{U}$.: there exists a constant $\mu_{\mathrm{e}}>0$ such that
    $\mathrm{dist}(x,S)\leq\mu_{\mathrm{e}}\cdot\mathrm{dist}(0,\hat{\partial}f(x)),\quad\forall x\in[f\leq f^{\star}+\nu].$ (EB)
  4. 4.
    Polyak-Łojasiewicz (PL) inequality [^3] To be consistent with the smooth case in [karimi2016linear], we call the property PL Polyak-Łojasiewicz, which is usually used for smooth functions. The property PL is actually a special case of the Kurdyka-Łojasiewicz inequality $\varphi^{\prime}(f(x)-f^{\star})\mathrm{dist}(0,\hat{\partial}f(x))\geq 1$ with $\varphi(s)=cs^{1/2}$ and $c>0$.: there exists a constant $\mu_{\mathrm{p}}>0$ such that
    $2\mu_{\mathrm{p}}\cdot(f(x)-f^{\star})\leq\mathrm{dist}^{2}(0,\hat{\partial}f(x)),\quad\forall x\in[f\leq f^{\star}+\nu].$ (PL)
  5. 5.
    Quadratic Growth (QG): there exists a constant $\mu_{\mathrm{q}}>0$ such that
    $\frac{\mu_{\mathrm{q}}}{2}\cdot\mathrm{dist}^{2}(x,S)\leq f(x)-f^{\star},\quad\forall x\in[f\leq f^{\star}+\nu].$ (QG)
All the regularity conditions above are defined over a sublevel set $f\leq f^{\star}+\nu]$. If $v=+\infty$, then they are global. In particular, SC imposes a quadratic lower bound at every point in the sublevel set. On the other hand, RSI, EB and PL all require a certain growth of the minimal norm element in $\hat{\partial}f(x)$ when moving away from its solution set $S$ or optimal value $f^{\star}$. It is easy to see that RSI, EB and PL all imply that every stationary point $0\in\hat{\partial}f(x)$ in the sublevel set $f\leq f^{\star}+\nu]$ is a global minimum (but they do not imply the uniqueness of stationary points). Finally, QG shows that $f(x)$ grows at least quadratically when moving away from the solution set $S$. Unlike RSI, EB and PL, this QG alone allows suboptimal stationary points.
Our first technical result summarizes the relationships among the five regularity conditions.
Theorem 3.1.
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed $\rho$-weakly convex function with $f^{\star}=\min_{x\in\mathbb{R}^{n}}f(x)$ and $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$. Suppose $S\neq\emptyset$ and let $\nu>0$ be the same constant throughout SC, RSI, EB, PL, and QG. The following relationship holds
[…]
Furthermore, if the QG coefficient satisfies $\mu_{\mathrm{q}}>\rho$ (including the case that $f$ is convex), then the following equivalence holds
[…]
Table 1: Constant conversion for RSI, EB, PL and QG, under the condition $\mu_{\mathrm{q}}>\rho$ in Theorem 3.1. Each row represents the corresponding coefficients for other conditions while holding one condition constant. For example, the second row shows the coefficients for EB, PL, and QG assuming that RSI holds with a positive constant $\mu_{\mathrm{r}}$.
RSI EB PL QG
RSI $\mu_{\mathrm{r}}$ $\frac{1}{\mu_{\mathrm{r}}},+\infty)$ $(0,\frac{\mu_{\mathrm{r}}^{2}}{2\mu_{\mathrm{r}}+\rho}]$ $(0,\frac{\mu_{\mathrm{r}}}{2}]$
EB $(0,\frac{1-2\mu_{\mathrm{e}}\rho}{4\mu_{\mathrm{e}}}]$ $\mu_{\mathrm{e}}$ $(0,\frac{1}{2\mu_{\mathrm{e}}+\rho\mu_{\mathrm{e}}^{2}}]$ $(0,\frac{1}{2\mu_{\mathrm{e}}}]$
PL $(0,\frac{\mu_{\mathrm{p}}-2\rho}{4}]$ $\frac{1}{\mu_{\mathrm{p}}},+\infty)$ $\mu_{\mathrm{p}}$ $(0,\frac{\mu_{\mathrm{p}}}{2}]$
QG $(0,\frac{\mu_{\mathrm{q}}-\rho}{2}]$ $\frac{2}{\mu_{\mathrm{q}}-\rho},+\infty)$ $(0,\frac{(\mu_{\mathrm{q}}-\rho)^{2}}{4\mu_{\mathrm{q}}}]$ $\mu_{\mathrm{q}}$
Note that if a function $f$ is convex and satisfies QG with $\mu_{\mathrm{q}}>0$, then 8 holds naturally as $f$ is naturally 0-weakly convex. Theorem 3.1 includes [karimi2016linear, Theorem 2] as a special case, in which only $L$-smooth functions are considered. It is easy to see that all $L$-smooth functions are also $L$-weakly convex. Even for differentiable functions, Theorem 3.1 is more general than [karimi2016linear, Theorem 2] in the sense that 1) we require no Lipschitz constant $L$ for gradients to ensure the equivalency among EB, PL, and QG in the convex case; 2) the condition $\mu_{\mathrm{q}}>\rho$ is new and does not mean that $f$ is convex. We will provide further illustrative examples in Section 3.2. Under the condition $\mu_{\mathrm{q}}>\rho$, we can further infer the coefficients for different regularity conditions from each other; the details are listed in Table 1.
The proof of Theorem 3.1 relies heavily on the notion of slope in Definition 1, Ekeland’s variational principle in Theorem 2.2, and the technical result in Lemma 2.2. We will present the proof details in Section 3.3. Alternative proofs based on subgradient flows [bolte2017error] are also possible. Indeed, one key step in the proof of [karimi2016linear, Theorem 2] is based on gradient flows for smooth functions, which is a special case of subgradient flows for nonsmooth cases.
Remark 1.
The regularity conditions EB, QG and PL have been discussed for different function classes in the literature. For the smooth case, we refer to [karimi2016linear, Theorem 2] for a nice summary; also see [rebjock2023fast, guille2021study] for related discussions. For nonsmooth convex functions, the equivalence between EB and QG has been recognized in [drusvyatskiy2018error, Theorem 3.3] and [artacho2008characterization, Theorem 3.3], and the equivalence between PL and QG is established in [bolte2017error, Theorem 5]. Thus, EB, PL, and QG are equivalent for the class of nonsmooth convex functions [ye2021variational, Proposition 2]; see also [zhu2023unified] for a recent discussion. Our Theorem 3.1 extends these results to $\rho$-weakly convex functions. The most closely related work is [drusvyatskiy2021nonsmooth] which focuses on nonsmooth optimization using Taylor-like models. Indeed, we specialize the proof of [drusvyatskiy2021nonsmooth, Theorem 3.7, Proposition 3.8, and Corollary 5.7] into our setting and prove the relationship in 8 and 7 directly using the slope technique. We note that the implication from QG to EB/PL is not true in general. Yet, with the condition $\mu_{\mathrm{q}}>\rho$ in Theorem 3.1, all four regularity conditions RSI, EB, PL and QG are equivalent. $\square$
3.2 Examples
In principle, all the five properties in Theorem 3.1 are generalizations of quadratic functions to non-quadratic, nonconvex, and even nonsmooth cases. For illustration, let us first consider the simplest quadratic function $f(x)=x^{2}$, which is convex and differentiable. It is clear that $\hat{\partial}f(x)=\{2x\}$ and $S=\{0\}$. It is also immediate to verify that SC holds with $0<\mu_{\mathrm{s}}\leq 2$, RSI holds with $0<\mu_{\mathrm{r}}\leq 2$, EB holds with $\mu_{\mathrm{e}}\geq 1/2$, PL holds with $0<\mu_{\mathrm{p}}\leq 2$, and QG holds with $0<\mu_{\mathrm{q}}\leq 2$. Consider another simple convex function $f(x)=x^{2},\text{if}\penalty 10000\ |x|\leq 1,$ and $f(x)=\frac{1}{2}x^{4}+\frac{1}{2}$ otherwise. All the five properties hold for this function, but it is not $L$-smooth globally.
Let us move away from convex functions, and consider $f(x)=x^{2}+6\sin^{2}(x)$. It is clear that this function satisfies QG globally, however, there exist suboptimal stationary points and consequently EB and PL do not hold globally. Thus, in this case (with $\nu=+\infty$), the relationship 7 is strict, and QG is more general than the other conditions.
Finally, we consider a $\rho$-weakly convex function with a QG constant satisfying $\mu_{\mathrm{q}}>\rho$ in the following example. Theorem 3.1 confirms that this function will also satisfy RSI, PL, and EB.
Example 1.
Consider the function
[…]
It can be verified that the function is not convex (due to the part 1<x<0.5-1<x<-0.5) but $2$-weakly convex with a unique minimizer $x^{\star}=-1$ and the minimum value $f^{\star}=0$. See Figure 1 for an illustration of these properties. The Fréchet subdifferential of this function can be calculated as
[…]
The length of the minimal element is then
[…]
As expected from Theorem 3.1, we can see that QG holds for $0<\mu_{\mathrm{q}}\leq 6$, EB holds for $\mu_{\mathrm{e}}\geq 1/2$, and PL holds for $0<\mu_{\mathrm{p}}\leq 2/3$. These can also be observed in Figure 1.
[…]
3.3 Proof of Theorem 3.1
The directions SC $\Rightarrow$ RSI, RSI $\Rightarrow$ EB, EB $\Rightarrow$ PL, and QG with $\mu_{\mathrm{q}}>\rho$ $\Rightarrow$ RSI follow easily from some elementary algebras. The directions PL $\Rightarrow$ EB and EB $\Rightarrow$ PL require the equivalence in Lemma 2.1. Most directions are not difficult to prove, except that the direction PL $\Rightarrow$ EB requires some sophisticated arguments involving Lemmas 2.2 and 2.2. To start the proof, we first recall that $S$ denotes the set of minimizers of $f$ and $f^{\star}=\min_{x\in\mathbb{R}^{n}}f(x)$ and we assume $S$ is nonempty.
  • SC $\Rightarrow$ RSI: Suppose $f$ satisfies SC with constant $\mu_{\mathrm{s}}>0$. For any $x\in[f\leq f^{\star}+\nu]$ and $g\in\hat{\partial}f(x)$, we let $\hat{x}\in\Pi_{S}(x)$. It follows that
[…]
    $\qquad\qquad\qquad\;\leq\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle$,
    where the last inequality comes from the fact that $f(\hat{x})-f(x)\leq 0$.
  • RSI $\Rightarrow$ EB: Suppose $f$ satisfies RSI with constant $\mu_{\mathrm{r}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $g$ be the minimal norm element in $\hat{\partial}f(x)$, and $\hat{x}\in\Pi_{S}(x)$. Then, by definition of RSI, we have
    $\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle\geq\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S).$
    Applying Cauchy-Schwarz on the left side yields
    $\mathrm{dist}(0,\hat{\partial}f(x))\geq\mu_{\mathrm{r}}\cdot\mathrm{dist}(x,S).$
  • EB $\Rightarrow$ PL: Suppose $f$ satisfies EB with constant $\mu_{\mathrm{e}}>0$. Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g$ be the minimal norm element in $\hat{\partial}f(x)$. From the subdifferential property for $\rho$-weakly convex function (i.e. 29 in Appendix A), we have
    $f^{\star}\geq f(x)+\mathopen{}\mathclose{{\left\langle g,\hat{x}-x}}\right\rangle-\frac{\rho}{2}\|\hat{x}-x\|^{2}.$
    Using the triangle inequality, it follows that
    $f(x)-f^{\star}$ $\leq\mathrm{dist}(0,\hat{\partial}f(x))\mathrm{dist}(x,S)+\frac{\rho}{2}\mathrm{dist}^{2}(x,S)$
    $\leq\mu_{\mathrm{e}}\cdot\mathrm{dist}^{2}(0,\hat{\partial}f(x))+\frac{\rho\mu_{\mathrm{e}}^{2}}{2}\mathrm{dist}^{2}(0,\hat{\partial}f(x))$
    $=\mathopen{}\mathclose{{\left(\frac{2\mu_{\mathrm{e}}+\rho\mu_{\mathrm{e}}^{2}}{2}}}\right)\mathrm{dist}^{2}(0,\hat{\partial}f(x)).$
    Dividing both sides by $(2\mu_{\mathrm{e}}+\rho\mu_{\mathrm{e}}^{2})/2$ gets the desired constant.
  • PL $\Rightarrow$ EB: Suppose $f$ satisfies PL with constant $\mu_{\mathrm{p}}>0$. As $f$ is $\rho$-weakly convex, applying Lemma 2.1 on PL shows that
    $|\nabla f|^{2}(x)\geq 2\mu_{\mathrm{p}}\cdot(f(x)-f^{\star}),\quad\forall x\in[f\leq f^{\star}+\nu].$ (9)
    Define the function $g$ as $x\mapsto g(x)=(f(x)-f^{\star})^{1/2}$. We claim that
    $|\nabla g|(x)\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\;\forall x\in[0<g\leq\sqrt{\nu}].$ (10)
    With 10, we can establish the implication PL $\Rightarrow$ EB as follows. Fix a point $x\in[0<g\leq\sqrt{\nu}]$ and choose $K>g(x)\sqrt{\frac{2}{\mu_{\mathrm{p}}}}$. We have $g(x)<K\sqrt{\frac{\mu_{\mathrm{p}}}{2}}$ and
    $|\nabla g|(u)$ $\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\quad\forall u\in[0<g\leq g(x)],\|x-u\|\leq K.$
    Choosing $r=\sqrt{\frac{\mu_{\mathrm{p}}}{2}},\alpha=0$ in Lemma 2.2, we have the following bounds
    $\mathrm{dist}(x,S)$ $=\mathrm{dist}(x,[g\leq 0])$
    $\leq\sqrt{\frac{2}{\mu_{\mathrm{p}}}}g(x)=\sqrt{\frac{2}{\mu_{\mathrm{p}}}}(f(x)-f^{\star})^{1/2}$
    $\leq\frac{1}{\mu_{\mathrm{p}}}|\nabla f|(x)=\frac{1}{\mu_{\mathrm{p}}}\mathrm{dist}(0,\hat{\partial}f(x)),\forall x\in[f\leq f^{\star}+\nu],$
    where the first inequality is from Lemma 2.2, the second inequality comes from 9, and the third equality applies Lemma 2.1. This shows the implication PL $\Rightarrow$ EB.
    Proof of the claim 10: Fix any $\bar{x}\in[0<g\leq\sqrt{\nu}]$. We shall show
    $|\nabla g|(\bar{x})\geq\sqrt{\frac{\mu_{\mathrm{p}}}{2}}.$ (11)
    Note that if $|\nabla g|(\bar{x})=+\infty$, then (11) holds trivially. Hence, we only prove (11) for the case $|\nabla g|(\bar{x})<+\infty$. By the definition of the slope $|\nabla f|(\bar{x})=\lim\sup_{y\rightarrow\bar{x}}\frac{(f(\bar{x})-f(y))^{+}}{\lVert y-\bar{x}\rVert}$ and by our choice of $\bar{x}$ satisfying $|\nabla f|(\bar{x})\geq 2\mu_{p}(f(\bar{x})-f^{\star})>0$, we claim that there exists a sequence $\{y_{i}\}$ satisfying the following conditions:
    $\{y_{i}\}\rightarrow\bar{x},\quad\{f(y_{i})\}\uparrow f(\bar{x}),\quad|\nabla f|(\bar{x})=\lim_{i\to+\infty}\frac{f(\bar{x})-f(y_{i})}{\lVert y_{i}-\bar{x}\rVert},$ (12)
    where $\{f(y_{i})\}\uparrow f(\bar{x})$ means $f(y_{i})\leq f(\bar{x})$ for all $i$ and $\{f(y_{i})\}\rightarrow f(\bar{x})$. If 12 holds, then (11) can be readily established using basic calculus. Indeed, using the Taylor expansion of the function $x\mapsto\sqrt{x}$ near $f(\bar{x})-f^{\star}$, we have
    $|\nabla g|(\bar{x})$ $=\limsup_{y\to\bar{x}}\frac{\mathopen{}\mathclose{{\left((f(\bar{x})-f^{\star})^{\frac{1}{2}}-(f(y)-f^{*})^{\frac{1}{2}}}}\right)^{+}}{\|y-\bar{x}\|}$
    $\geq\limsup_{i\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(y_{i})-f^{*})^{\frac{1}{2}}}{\lVert y_{i}-\bar{x}\rVert}$
    $=\limsup_{i\rightarrow+\infty}\frac{\frac{1}{2(f(\bar{x})-f^{*})^{\frac{1}{2}}}(f(\bar{x})-f(y_{i}))+o(|f(y_{i})-f(\bar{x})|)}{\lVert y_{i}-\bar{x}\rVert}$
    $\overset{(a)}{=}\frac{|\nabla f|(\bar{x})}{2(f(\bar{x})-f^{*})^{\frac{1}{2}}}$
    $\overset{(b)}{\geq}\sqrt{\frac{\mu_{\mathrm{p}}}{2}},$
    where $(a)$ is due to (12), and $(b)$ applies (9). It remains to prove the claim 12.
    Proof of the claim 12: The conditions $\{y_{i}\}\rightarrow\bar{x}$, $f(y_{i})\leq f(\bar{x})$ for all $i$, and $|\nabla f|(\bar{x})=\lim_{i\to+\infty}\frac{f(\bar{x})-f(y_{i})}{\lVert y_{i}-\bar{x}\rVert}$ follow from the definition of $|\nabla f|(\bar{x})$ and the assumption $|\nabla f|(\bar{x})>0$. Below, we further establish $\{f(y_{i})\}\uparrow f(\bar{x})$. Suppose for contradiction that the sequence $\{f(y_{i})\}$ does not converge to $f(\bar{x})$. Then, let $\epsilon=\frac{1}{2}(f(\bar{x})-f^{\star})>0$, there exists a subsequence $\{f(y_{i_{k}})\}_{k}$ such that $\epsilon\leq f(\bar{x})-f(y_{i_{k}})$. However, this subsequence will result in
    $|\nabla g(\bar{x})|$ $\geq\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(y_{i_{k}})-f^{*})^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $\geq\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(f(\bar{x})-f^{*}-\epsilon)^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $=\limsup_{k\rightarrow+\infty}\frac{(f(\bar{x})-f^{*})^{\frac{1}{2}}-(\frac{1}{2}(f(\bar{x})-f^{*}))^{\frac{1}{2}}}{\lVert y_{i_{k}}-\bar{x}\rVert}$
    $=+\infty,$
    violating the assumption $|\nabla g|(\bar{x})<+\infty$. This proves the claim 12.
  • EB $\Rightarrow$ QG: Suppose $f$ satisfies EB with $\mu_{\mathrm{e}}>0$. Fixing any $x\in[f^{\star}<f\leq f^{\star}+\nu]$, we have $x\in[f^{\star}<f\leq f^{\star}+\epsilon]$ with $\epsilon=f(x)-f^{\star}$. Choosing $\rho=\sqrt{\epsilon/\mu_{\mathrm{e}}}$, Theorems 2.2 and 6 ensure that there exists a $y$ such that $y\in[f\leq f^{\star}+\epsilon]$ and $|\nabla f|(y)\leq\rho$. Thus,
    $\mu_{\mathrm{e}}\rho\geq\mu_{\mathrm{e}}\cdot|\nabla f|(y)$ $\overset{(a)}{\geq}\mathrm{dist}(y,S)$
    $\overset{(b)}{\geq}\mathrm{dist}(x,S)-\mathrm{dist}(x,y)$
    $\overset{(c)}{\geq}\mathrm{dist}(x,S)-\epsilon/\rho,$
    where $(a)$ uses Lemma 2.1 (i.e., we have $|\nabla f|(y)=\mathrm{dist}(0,\hat{\partial}f(y))$ for weakly convex functions) and EB, $(b)$ applies the triangle inequality for a point to a set, and $(c)$ comes from 5a. Substituting $\epsilon=f(x)-f^{\star}$ and $\rho=\sqrt{\frac{\epsilon}{\mu_{\mathrm{e}}}}$ and rearranging the above inequality yields
    $2\sqrt{\mu_{\mathrm{e}}}(f(x)-f^{\star})^{1/2}\geq\mathrm{dist}(x,S).$
    Squaring both sides completes the proof of EB $\Rightarrow$ QG.
Summarizing the relationship above leads to the first part of Theorem 3.1 in 7. We now prove the second part in 8.
  • QG with $\mu_{\mathrm{q}}>\rho$ $\Rightarrow$ RSI: Let $x\in[f\leq f^{\star}+\nu]$, $\hat{x}\in\Pi_{S}(x)$, and $g\in\hat{\partial}f(x)$. From the assumption of QG and the property of subgradient of weakly convex function in 29, we have
    $\frac{\mu_{\mathrm{q}}}{2}\cdot\mathrm{dist}^{2}(x,S)\leq f(x)-f^{\star}$ $\leq\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle+\frac{\rho}{2}\mathrm{dist}^{2}(x,S).$
    Rearranging terms yields
    $\mathopen{}\mathclose{{\left(\frac{\mu_{\mathrm{q}}-\rho}{2}}}\right)\cdot\mathrm{dist}^{2}(x,S)\leq\mathopen{}\mathclose{{\left\langle g,x-\hat{x}}}\right\rangle$.
This completes our proof of Theorem 3.1.
4 Proximal point method for convex optimization
In this section, we will utilize the regularity conditions in Section 3 to derive fast linear convergence guarantees of the classical proximal point method (PPM) [rockafellar1976augmented] for convex (potentially nonsmooth) optimization and generalize it to weakly convex settings. PPM is a conceptually simple algorithm, which has been historically used for guiding algorithm design and analysis, such as proximal bundle methods [lemarechal1981bundle] and augmented Lagrangian methods [rockafellar1976augmented]. It has recently found increasing applications in modern machine learning; see [drusvyatskiy2017proximal].
4.1 Proximal point method
Consider the optimization problem
$f^{\star}=\min_{x}\;f(x),\vskip-2.84526pt$ (13)
where $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is a proper closed convex function. Note that 13 is also an abstract model for constrained optimization, since given a closed convex set $X$, we can define $\bar{f}(x)=f(x)$ if $x\in X$, and $\bar{f}(x)=+\infty$ otherwise. Let $S=\operatorname*{\arg\!\min}_{x}\,f(x)$. We define the proximal mapping as
$\text{prox}_{\alpha f}(x_{k}):=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}\;f(x)+\frac{1}{2\alpha}\mathopen{}\mathclose{{\left\|x-x_{k}}}\right\|^{2},$ (14)
where $\alpha>0$. Starting with any initial point $x_{0}$, the PPM generates a sequence of points as follows
$x_{k+1}=\mathrm{prox}_{c_{k}f}(x_{k}),\quad k=0,1,2,\ldots$ (15)
where $\{c_{k}\}_{k\geq 0}$ is a sequence of positive real numbers. The quadratic term in 14 makes the objective function strongly convex and always admits a unique solution. The iterates 15 are thus well-defined.
The convergence of PPM 15 for (nonsmooth) convex optimization has been studied since the 1970s [rockafellar1976monotone]. The sublinear convergence is relatively easy to establish, and many different assumptions exist for linear convergences of 15; see [rockafellar1976monotone, luque1984asymptotic, leventhal2009metric, cui2016asymptotic, drusvyatskiy2018error]. However, as we will highlight later, some assumptions are restrictive and the corresponding proofs are sophisticated and nontransparent. We aim to provide simple proofs under the general regularity conditions in Section 3.
4.2 (Sub)linear convergences of PPM
Under a very general setup, the PPM 15 converges at a sublinear rate for cost value gaps, and the iterates converge asymptotically, as summarized in Theorem 4.1. This result is classical [guler1991convergence, Theorem 2.1], and a new bound with an improved constant 4 is also available in [taylor2017exact, Theorem 4.1] using the performance estimation technique.
Theorem 4.1 (Sublinear convergence [guler1991convergence, Theorem 2.1]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Then, the iterates 15 with a positive sequence $\{c_{k}\}_{k\geq 0}$ satisfy
$f(x_{k})-f^{\star}\leq\mathrm{dist}^{2}(x_{0},S)/(2\textstyle\sum_{t=0}^{k-1}c_{t}).$ (16)
If we further have $\lim_{k\to\infty}\sum_{t=0}^{k-1}c_{t}=\infty$, then the iterates converge to an optimal solution $\bar{x}$ asymptotically, i.e., $\lim_{k\to\infty}x_{k}=\bar{x}$, where $\bar{x}\in S$.
The proof of 16 is immediate from a telescope sum via the following one-step improvement:
$2c_{k}(f(x_{k+1})-f(x^{\star}))\leq\|x_{k}-x^{\star}\|^{2}-\|x_{k+1}-x^{\star}\|^{2},\quad\forall c_{k}>0,x^{\star}\in S.$ (17)
This fact is not difficult to establish. Note that choosing any constant step size $c_{k}=c>0$ in 16 directly implies the common sublinear rate $\mathcal{O}(1/k)$. In Theorem 4.1, $f$ does not need to be $L$-smooth, and it can also be non-differentiable. Thus, the guarantees in Theorem 4.1 are much stronger than those by (sub)gradient methods. This is because the proximal mapping 14 is a stronger oracle than simple (sub)gradient updates.
Similar to GD in Section 2, when the function $f$ satisfies additional regularity conditions, the PPM enjoys linear convergence. Our next main technical result establishes linear convergences of the PPM under the common regularity conditions in Theorem 3.1.
Theorem 4.2 (Linear convergence).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x), SS\neq\emptyset$, and $\nu>0$. Suppose $f$ satisfies PL (or EB, RSI, QG) over the sublevel set $f\leq f^{\star}+\nu]$. Then, for all $k\geq k_{0}=\frac{\mathrm{dist}^{2}(x_{0},S)}{2\nu\inf_{k\geq 0}c_{k}}$ steps, the iterates 15 with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero enjoy linear convergence rates, i.e.,
$f(x_{k+1})-f^{\star}$ $\leq\omega_{k}\cdot(f(x_{k})-f^{\star}),$ (18a)
$\mathrm{dist}(x_{k+1},S)$ $\leq\theta_{k}\cdot\mathrm{dist}(x_{k},S),$ (18b)
where the constants are
$\omega_{k}=\frac{1}{1+c_{k}\mu_{\mathrm{p}}}<1,\quad\theta_{k}=\min\mathopen{}\mathclose{{\left\{\frac{1}{\sqrt{c_{k}\mu_{\mathrm{q}}+1}},\frac{1}{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}}}\right\}<1.$
Proof.
The sublinear convergence in Theorem 4.1 ensures that the iterate $x_{k}$ reaches $f\leq f^{\star}+\nu]$ after at most $k_{0}$ iterations. Once $x_{k}$ is within $f\leq f^{\star}+\nu]$, all the properties EB, PL, RSI, and QG are equivalent by Theorem 3.1, and the next point $x_{k+1}$ also satisfies $x_{k+1}\in[f\leq f^{\star}+\nu]$ as $f(x_{k+1})\leq f(x_{k+1})+\frac{1}{2c_{k}}\|x_{k+1}-x_{k}\|^{2}\leq f(x_{k})$ by the update rule 15. For the analysis below, we assume $x_{k}\in[f\leq f^{\star}+\nu]$.
We next show that PL gives a simple proof of 18a, and QG together with EB leads to a clean proof of 18b. Recall that the optimality condition of 15 directly implies
$(x_{k+1}-x_{k})/c_{k}\in\partial f(x_{k+1}).$ (19)
Then, the following inequalities hold
$f(x_{k})-f(x_{k+1})$ $\overset{(a)}{\geq}\|x_{k+1}-x_{k}\|^{2}/(2c_{k})$ (20)
$\overset{(b)}{\geq}\frac{c_{k}}{2}\mathrm{dist}^{2}(0,\partial f(x_{k+1}))\overset{(c)}{\geq}c_{k}\mu_{\mathrm{p}}(f(x_{k+1})-f^{\star}),$
where $(a)$ applies the fact that $x_{k}$ is a suboptimal solution to 15, $(b)$ comes from the optimality 19, and $(c)$ applies PL. Re-arranging and subtracting $f^{*}$ from both sides of 20 lead to the desired linear convergence in 18a.
We next use QG to prove 18b with coefficient $\theta_{k}\leq 1/\sqrt{c_{k}\mu_{\mathrm{q}}+1}$. By definition, we have $f(\Pi_{S}(x_{k}))=f^{\star}$ and $\|\Pi_{S}(x_{k})-x_{k}\|^{2}=\mathrm{dist}^{2}(x_{k},S)$. Since $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $1/c_{k}$ strongly convex, its first-order lower bound at $x_{k+1}$ is
$f^{\star}+\frac{1}{2c_{k}}\|\Pi_{S}(x_{k})-x_{k}\|^{2}=f(\Pi_{S}(x_{k}))+\frac{1}{2c_{k}}\mathrm{dist}^{2}(x_{k},S)$ (21)
[…]
where we also applied the fact that $x_{k+1}$ minimizes 15 so 0 is a subgradient. From 21, we drop the positive term $\|x_{k+1}-x_{k}\|^{2}$ and use the fact that $\|\Pi_{S}(x_{k})-x_{k+1}\|\geq\mathrm{dist}(x_{k+1},S)$, leading to
$f^{\star}-f(x_{k+1})+\mathrm{dist}^{2}(x_{k},S)/(2c_{k})\geq\mathrm{dist}^{2}(x_{k+1},S)/(2c_{k}).$
Combining this inequality with QG and simple re-arranging terms leads to the desired linear rate
$\mathrm{dist}^{2}(x_{k+1},S)\leq 1/\sqrt{c_{k}\mu_{\mathrm{q}}+1}\cdot\mathrm{dist}^{2}(x_{k},S).$
Simple arguments based on EB can establish 18b with coefficient $\theta_{k}\leq 1/(\mu_{\mathrm{e}}^{2}/c_{k}^{2}+1)^{1/2}$. We provide some details in Section C.1. ∎
Two nice features of Theorem 4.2 are 1) the simplicity of its proofs and 2) the generality of its conditions. Indeed, the proof of 18a is simple via PL, and the proof of 18b is clean via QG and EB, which are simpler than typical proofs. In addition, the regularity conditions are weaker than [rockafellar1976monotone] and [luque1984asymptotic]. Linear convergence for $\mathrm{dist}(x_{k},S)$ was first established in [rockafellar1976monotone] with one restrictive assumption: the inverse of the subdifferential $(\partial f)^{-1}$ is locally Lipschitz at 0, which requires a unique optimal solution, i.e., $S$ is a singleton. The uniqueness assumption is lifted in [luque1984asymptotic], which even allows an unbounded solution set. More recently, this assumption is further relaxed to $\partial f$ being metrically subregular in [cui2016asymptotic] and [leventhal2009metric]. It is known that for convex functions, $\partial f$ is metrically subregular if and only if $f$ satisfies quadratic growth (cf. Theorem 3.1). Indeed, our proof in Theorem 4.2 is based on quadratic growth, which is a more intuitive geometrical property. Our main idea in the proof above is partially inspired by a recent result for the linear convergence of the spectral bundle method in [ding2023revisiting, liao2023overview].
We note that in the proof of Theorem 4.2, once the iterate is within the PL/QG region, the proof idea follows simply on the existence of the optimality condition 19. Building upon this observation, we extend Theorem 4.2 to the class of weakly convex functions when a proper initialization is given.
Theorem 4.3 (Extension to weakly convex functions).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed $\rho$-weakly convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies QG with $\mu_{\mathrm{q}}>\rho$ over the sublevel set $f\leq f^{\star}+\nu]$ with $\nu>0$. Let $\beta=\mu_{\mathrm{q}}-\rho>0$. The iterates generated by the PPM 15 with any positive sequence $\{c_{k}\}_{k\geq 0}$ satisfying $1/c_{k}>\rho$ for all $k\geq 0$ and an initialization $x_{0}\in[f\leq f^{\star}+\nu]$ satisfy
$f(x_{k+1})-f^{\star}$ $\leq\omega_{k}\cdot(f(x_{k})-f^{\star}),$
$\mathrm{dist}(x_{k+1},S)$ $\leq\theta_{k}\cdot\mathrm{dist}(x_{k},S),$
where the constants are
$\omega_{k}=\frac{1}{1+\mu_{\mathrm{p}}c_{k}}<1\penalty 10000\ \text{and}\penalty 10000\ \theta_{k}=\min\mathopen{}\mathclose{{\left\{\frac{1}{\sqrt{c_{k}\beta+1}},\frac{1}{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}}}\right\}<1,$
with $\mu_{\mathrm{p}}$ and $\mu_{\mathrm{e}}$ being any valid constants for PL and EB from Table 1, respectively.
Proof.
We first note that choosing $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho,\forall k\geq 0$, guarantees the existence of the optimality condition $(x_{k+1}-x_{k})/c_{k}\in\hat{\partial}f(x_{k+1})$ similar to 19. With the existence of the optimality condition, the proof follows similar arguments in the proof of Theorem 4.2. Specifically, the proof of the linear decrease in the cost value gap is identical to that in Theorem 4.2. As $\mu_{\mathrm{q}}>\rho$, EB, RSI, and PL hold by Theorem 3.1. For the proof of the linear decrease in the distance, the coefficient $(c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1)^{-1/2}$ follows the same argument as in the convex case since $c_{k}$ is chosen such that $1/c_{k}>\rho$ so the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is a convex function plus a quadratic term and the proximal mapping satisfied the key inequality in Lemma C.1 in Section C.1; For the coefficient $(c_{k}\beta+1)^{-1/2}$, the argument is essentially the same as the convex case with a small modification that the function $f+\frac{1}{2c_{k}}\|\cdot-x_{k}\|^{2}$ is $(\frac{1}{c_{k}}-\rho)$-strongly convex. ∎
Note that Theorem 4.3 requires the initial point to be within the sublevel set that satisfies QG and $\mu_{\mathrm{q}}>\rho$. If the function $f$ satisfies QG and $\mu_{\mathrm{q}}>\rho$ globally (i.e., $\nu=+\infty$), the above linear convergence holds with any initial condition. Let us illustrate the linear convergence for weakly convex function in the following example.
Example 2 (PPM for a weakly convex function).
Consider the 2-weakly convex function $f$ in Example 1. It satisfies QG and $6=\mu_{\mathrm{q}}>\rho=2$ globally. Therefore, any initial point $x_{0}\in\mathbb{R}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ such that $1/c_{k}>\rho$ for all $k\geq 0$ guarantees the linear convergence. Let $c_{k}=1/6$ for all $k\geq 0$. For any $x_{k}\in\mathbb{R}$, from a simple calculation, the PPM update 15 follows the formula
$x_{k+1}=\operatorname*{\arg\!\min}_{v\in\{v_{1},v_{2}\}}f(v)+3(v-x_{k})^{2},$
where
$v_{1}=\begin{cases}-0.5,&\penalty 10000\ \text{if}\penalty 10000\ \frac{3}{2}x_{k}\geq-0.5\\ \frac{3}{2}x_{k},&\penalty 10000\ \text{if}\penalty 10000\ -1<\frac{3}{2}x_{k}<-0.5\\ -1,&\penalty 10000\ \text{if}\penalty 10000\ \frac{3}{2}x_{k}\leq-1\end{cases}$
and
$v_{2}=\begin{cases}\operatorname*{\arg\!\min}_{x\in\{-1,-0.5\}}f(x)+3(x-x_{k})^{2},&\penalty 10000\ \text{if}\penalty 10000\ -1<\frac{x_{k}-1}{2}<-0.5\\ \frac{x_{k}-1}{2},&\penalty 10000\ \text{otherwise}.\end{cases}$
We present the numerical result with the initial point $x_{0}=-2$ in Figure 2. It shows the linear convergence in both the distance to the solution set and the cost value gap.
Figure 2: Linear convergences of the distance to the solution set (left) and cost value gaps (right) for a $2$-weakly convex function in Example 2.
5 Inexact proximal point method (iPPM) and its convergence
In Section 4, each subproblem 14 is solved exactly. This may not be practical since one still needs an iterative solver to solve 14, where stopping criteria naturally introduce errors. We here discuss an inexact version of PPM (iPPM, [rockafellar1976monotone]) where the subproblem 15 is solved inexactly. The regularity conditions in Theorem 3.1 also allow us to establish linear convergences of iPPM.
5.1 iPPM and stopping criteria
We replace the exact update 15 with an inexact update
$x_{k+1}\approx\text{prox}_{c_{k}f}(x_{k}).$ (23)
Two classical criteria suggested in Rockafellar’s seminal work [rockafellar1976monotone] are
$\|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|$ $\leq\epsilon_{k},\quad\textstyle\sum_{k=0}^{\infty}\epsilon_{k}<\infty,$ (A)
$\|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|$ $\leq\delta_{k}\|x_{k+1}-x_{k}\|,\quad\textstyle\sum_{k=0}^{\infty}\delta_{k}<\infty.$ (B)
The inexact update 23 with A or B is called iPPM. The two criteria are not implementable as the value of $\text{prox}_{c_{k}f}(x_{k})$ is unknown. As discussed in [rockafellar1976monotone, Proposition 3], two implementable alternatives that imply A and B, respectively, are
$\mathrm{dist}(0,H_{k}(x_{k+1}))$ $\leq\epsilon_{k}/c_{k},\quad\textstyle\sum_{k=0}^{\infty}\epsilon_{k}<\infty,$ ($\text{A}^{\prime}$)
$\mathrm{dist}(0,H_{k}(x_{k+1}))$ $\leq(\delta_{k}/c_{k})\|x_{k+1}-x_{k}\|,\quad\textstyle\sum_{k=0}^{\infty}\delta_{k}<\infty,$ ($\text{B}^{\prime}$)
where $H_{k}(x)=\partial f(x)+(x-x_{k})/c_{k}$ is the subdifferential of $f+\|\cdot-x_{k}\|^{2}/(2c_{k})$ at $x$ (since $f$ is convex by assumption). Note that A and B only require the inexact update $x_{k+1}$ to stay close enough to $\text{prox}_{c_{k}f}(x_{k})$ with respect to the Euclidean distance, but they do not require the inexact update $x_{k+1}$ to be within the domain of $f$, i.e., $f(x_{k+1})$ might be infinity. However, the stopping criteria ($\text{A}^{\prime}$) and ($\text{B}^{\prime}$) require that $x_{k+1}$ is in the domain of $f$.
5.2 (Sub)linear convergence of iPPM
The seminal work by Rockafellar [rockafellar1976monotone] has established the asymptotic convergence of iterates for iPPM under a general setup. We state the results below, whose proof is technically involved.
Theorem 5.1 (Asymptotic convergence of iterates [rockafellar1976monotone, Theorem 1]).
Consider a proper closed convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by 23 under A with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, we have 1) the sequence $\{x_{k}\}_{k\geq 0}$ is bounded if and only if there exists a solution to $0\in\partial f(x)$, i.e., $S\neq\emptyset$; 2) if $S\neq\emptyset$, the whole sequence $\{x_{k}\}_{k\geq 0}$ converges to an optimal point $x_{\infty}\in S$, i.e., $\lim_{k\to\infty}x_{k}=x_{\infty}$.
We next establish a sublinear convergence of iPPM for cost value gaps. Our simple proof is based on the boundedness of the iterates from Theorem 5.1 and a recent idea in [lu2023unified, Theorem 3]. We provide its proof details in Appendix D.
Theorem 5.2 (Sublinear convergence of iPPM).
Let $f:\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. The iterates 23 under $\text{A}^{\prime}$ with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero converge to $x^{\star}\in S$ asymptotically, and the cost value gaps satisfy
$\textstyle\min_{j=0,\ldots,k}f(x_{j})-f^{\star}\leq(\mathrm{dist}^{2}(x_{0},S)+2D\sum_{j=0}^{k-1}\epsilon_{j})/(2\sum_{j=0}^{k-1}c_{j}),$
where $D$ is the diameter of iterates $\{x_{k}\}_{k\geq 0}$ which is bounded.
Note that Theorems 5.2 and 5.1 can be viewed as the convergence counterpart for iPPM of Theorem 4.1 with two major differences: 1) Theorem 5.2 deals with the best iterate, unlike the last iterate in Theorem 4.1 (the guarantee for the average $\bar{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x_{j}$ or weighted average $\tilde{x}_{k}=({\sum_{j=0}^{k-1}c_{j}x_{j+1}})/({\sum_{j=0}^{k-1}c_{j}})$ is also straightforward; see remark 3 in the appendix); 2) the convergence of cost values in Theorem 5.2 relies on the boundedness of iterates in Theorem 5.1 whose proof is technically involved [rockafellar1976monotone, Theorem 1], while the convergence of iterates of exact PPM is established from the sublinear convergence of cost values in Theorem 4.1.
Similar to Theorem 4.2, linear convergence of iPPM can also be deduced when popular regularity conditions are assumed. However, the definitions of SC, RSI, EB, PL, and QG in Section 3 are stated over a sublevel set $f\leq f^{\star}+\nu]$ with $\nu>0$, which is convenient to estimate the iteration number to enter a certain region, such as the constant $k_{0}$ in Theorem 4.2; but those definitions become too stringent in the analysis. Indeed, note that the proof in Theorem 4.2 requires the next point $x_{k+1}$ remains inside the QG or PL region. As the stopping criteria A and B allow infeasible points, it is possible that $x_{k+1}$ is infeasible (i.e., $f(x_{k+1})=+\infty$), making the same proof fail. However, the iterates are still approaching an optimal solution and the inequality $\mu_{\mathrm{q}}/2\cdot\mathrm{dist}(x_{k+1},S)\leq f(x_{k+1})-f^{\star}$ with any $\mu_{\mathrm{q}}>0$ still holds automatically.
To state the linear convergence result for iPPM, we here modify the definition of QG from a sublevel set to a neighborhood of an optimal solution. Precisely, we say $f$ satisfies QG at $x^{\star}\in S$, if there is a constant $\mu_{\mathrm{q}}>0$ and a neighborhood $\mathcal{U}\subseteq\mathbb{R}^{n}$ containing $x^{\star}$ such that
$\frac{\mu_{\mathrm{q}}}{2}\cdot\mathrm{dist}^{2}(x,S)\leq f(x)-f^{\star},\quad\forall x\in\mathcal{U}.$ (24)
Notice that we can also redefine RSI, EB, and PL using a neighborhood (similar to 24) and show the equivalence between them. However, we note that the neighborhood $\mathcal{U}$ for different regularity conditions may be different. For clarity, we present our final technical result that shows the linear convergence of iPPM under 24, which is the counterpart of Theorem 4.2.
Theorem 5.3 (Linear convergence of iPPM).
Let $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Suppose $f$ satisfies 24 at all $x^{\star}\in S$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence generated by iPPM 23 under A and B with a positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then, there exist a nonnegative $\theta_{k}<1$ and a large $\bar{k}>0$ such that
$\mathrm{dist}(x_{k+1},S)\leq\hat{\theta}_{k}\mathrm{dist}(x_{k},S),\;\forall k\geq\bar{k},\quad\text{where}\quad\hat{\theta}_{k}=\frac{\theta_{k}+2\delta_{k}}{1-\delta_{k}}<1.$
To prove Theorem 5.3, we first introduce an important inequality characterizing the one-step improvement, which was first established in [luque1984asymptotic, Equation 2.7]. We provide some details in Section D.2 for completeness.
Lemma 5.1 (Inexact one-step improvement [luque1984asymptotic, Equation 2.7]).
Let $f\!:\!\mathbb{R}^{n}\!\to\!\overline{\mathbb{R}}$ be a proper closed convex function with $S=\operatorname*{\arg\!\min}_{x\in\mathbb{R}^{n}}f(x)$ and $S\neq\emptyset$. Let $\{x_{k}\}_{k\geq 0}$ be any sequence by the iPPM 23 under B with the positive sequence $\{c_{k}\}_{k\geq 0}$ bounded away from zero. Then there exists a $\hat{k}\geq 0$ such that
$(1-\delta_{k})\mathrm{dist}(x_{k+1},S)\leq 2\delta_{k}\mathrm{dist}(x_{k},S)+\mathrm{dist}(\mathrm{prox}_{c_{k}f}(x_{k}),S),\;\forall k\geq\hat{k}$, (25)
where $\delta_{k}<1$.
Proof of Theorem 5.3: Thanks to A, the convergence of $\{x_{k}\}_{k\geq 0}$ is guaranteed by Theorem 5.1, i.e., the sequence $\{x_{k}\}_{k\geq 0}$ converges to an optimal solution $x^{\star}$. Let $\mu_{\mathrm{q}}$ and $\mathcal{U}$ be the positive constant and the neighborhood that satisfy
$\frac{\mu_{\mathrm{q}}}{2}\cdot\mathrm{dist}^{2}(x,S)\leq f(x)-f^{\star},\quad\forall x\in\mathcal{U}$ (26)
with $x^{\star}\in\mathcal{U}.$ Let $r$ be the largest radius such that $\mathbb{B}_{\mathbb{R}^{n}}(x^{\star},r)\subseteq\mathcal{U}$. Since $\{x_{k}\}_{k\geq 0}$ converges to $x^{\star}$ and stopping criterion A is implemented, there exist two integers $k_{1},k_{2}\geq 0$ such that
$\|x_{k}-x^{\star}\|<\frac{r}{2},$ $\;\forall k\geq k_{1},\penalty 10000\ \text{and}\penalty 10000\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|<\frac{r}{2},\;\forall k\geq k_{2}.$
Therefore, for all $k\geq\max\{k_{1},k_{2}\}$, we have
$\|\text{prox}_{c_{k}f}(x_{k})-x^{\star}\|\leq\|\text{prox}_{c_{k}f}(x_{k})-x_{k+1}\|+\|x_{k+1}-x^{\star}\|<r,$
which implies the point $\text{prox}_{c_{k}f}(x_{k})$ satisfies 26. The same analysis in Theorem 4.2 shows
$\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S)\leq\theta_{k}\mathrm{dist}(x_{k},S),\;\forall k\geq\max\{k_{1},k_{2}\},$ (27)
where $\theta_{k}=1/\sqrt{c_{k}\mu_{\mathrm{q}}+1}<1$. On the other hand, thanks to B and 5.1, there exists a $k_{3}>0$ such that
$(1-\delta_{k})\mathrm{dist}(x_{k+1},S)\leq 2\delta_{k}\mathrm{dist}(x_{k},S)+\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S),\;\forall k\geq k_{3},$ (28)
Consequently, combining 27 with 28 yields
$\mathrm{dist}(x_{k+1},S)\leq\frac{\theta_{k}+2\delta_{k}}{1-\delta_{k}}\mathrm{dist}(x_{k},S),\quad\forall k\geq\max\{k_{1},k_{2},k_{3}\}.$
Finally, as we know $\delta_{k}\to 0$ and $\theta_{k}<1,$ for all $k\geq\max\{k_{1},k_{2}\},$ there exists a $\bar{k}\geq\max\{k_{1},k_{2},k_{3}\}$ such that
$\frac{\theta_{k}+2\delta_{k}}{1-\delta_{k}}<1,\quad\forall k\geq\bar{k}.$
This completes the proof. ∎
In Theorem 5.3, criterion A serves to guarantee that the iterates can reach the neighborhood $\mathcal{U}$ in 24 (cf. the asymptotic convergence from Theorem 5.1). If 24 holds globally (i.e., $r=+\infty$), the same linear convergence result holds with B only. Note that the convergence proof in Theorem 5.3 is very modular. Indeed, thanks to the regularity conditions (i.e., 24), our proof appears simpler and less conservative than typical proofs in the literature; see the discussions after Theorem 4.2.
6 Applications
In this section, we present numerical experiments demonstrating the linear convergence of PPM. Throughout the experiment, we use the modeling package YALMIP [Lofberg2004] to formulate the subproblem 15 and call the conic solver MOSEK [mosek] to solve it. Our code is available at
https://github.com/soc-ucsd/PPM_examples
We consider three different applications of convex optimization in machine learning and signal processing: linear support vector machine (SVM) [zhang2015stochastic], lasso [tibshirani1996regression], and elastic-net [zou2005regularization], respectively. The problem formulations are
  • Linear SVM:
    $\min_{x\in\mathbb{R}^{n}}\;\frac{1}{n}\sum_{i=1}^{n}\max\{0,1-b_{i}(a_{i}^{\mathsf{T}}x)\}+\frac{\rho}{2}\|x\|^{2},$
    where $\rho>0,a_{i}\in\mathbb{R}^{d},b_{i}\in\{-1,1\},i=1,\ldots,n.$
  • Lasso ($\ell_{1}$-regularization):
    $\min_{x\in\mathbb{R}^{m}}\quad\frac{1}{2}\|y-Ax\|^{2}+\lambda\|x\|_{1},$
    where $A\in\mathbb{R}^{n\times m},y\in\mathbb{R}^{m},$ and $\lambda>0$.
  • Elastic-Net ($\ell_{1}-\ell_{2}^{2}$-regularization):
    $\min_{x\in\mathbb{R}^{m}}\quad\frac{1}{2}\|y-Ax\|^{2}+\lambda\|x\|_{1}+\frac{\mu}{2}\|x\|^{2},$
    where $A\in\mathbb{R}^{n\times m},y\in\mathbb{R}^{m},$ and $\lambda,\mu>0$.
For linear SVM, we consider three datasets (“a1a”, “Australian”, “fourclass”) from [chang2011libsvm]. The regularization term is set as $\rho=1.$ The step size parameter of the PPM is chosen as $c_{k}=1$ for all $k\geq 1$. For Lasso, we randomly generate $A\in\mathbb{R}^{n\times m}$ and $\hat{x}\in\mathbb{R}^{m}$ with $s$ elements being zero ($s<m$), then construct the vector $y=A\hat{x}$. We consider problems with three different sizes: $(n=10,m=40,s=5),(n=20,m=50,s=10),$ and $(n=30,m=60,s=15)$. The regularization terms are set as $\rho=10$. The step size parameter of the PPM is chosen as $c_{k}=0.16$ for all $k\geq 1$. For Elastic-Net, we consider the same data in Lasso with the regularization terms $\rho=10$ and $\mu=1.$ These problems satisfy the regularity conditions in Theorem 4.2. For each application, we run the PPM on three different data sets. The numerical results are shown in Figure 3, which confirms the fast linear convergence of the PPM.
Figure 3: Linear convergences of cost value gaps for linear SVM (left), lasso (middle), and elastic-net (right).
7 Conclusion
In this paper, we have established the relationship between different regularity conditions under the class of $\rho$-weakly convex functions (see Theorem 3.1). The result shows that, in the class of $\rho$-weakly convex functions, strong convexity implies restricted secant inequality, which in turn implies error bound. Meanwhile, error bound and Polyak-Łojasiewicz inequality are equivalent, and any of them implies quadratic growth. If a $\rho$-weakly convex function satisfies quadratic growth and the quadratic growth constant satisfies $\mu_{\mathrm{q}}>\rho$, then all these four properties: restricted secant inequality, error bound, Polyak-Łojasiewicz inequality, and quadratic growth, become equivalent. Unlike previous results in the literature, our result goes beyond smoothness and convexity, benefiting the analysis of various iterative algorithms.
We have also presented simple and modular proofs for the exact PPM for both convex and weakly convex functions which makes the analysis of the exact PPM more accessible (see Theorems 4.2 and 4.3). In the setting of inexact PPM, we first clarified the subtly in terms of inexactness and infeasibility, and then established the linear convergence of inexact PPM when the inexactness is controlled properly for convex optimization (see Theorem 5.3). The numerical experiments also validate our theoretical findings. We believe these results on (inexact) PPM will facilitate algorithm development in nonsmooth optimization, and we are particularly interested in further applications in large-scale conic optimization [zheng2021chordal, zheng2020chordal].
Appendix A Subdifferential characterization of weakly convex functions
For self-completeness, we review a useful subdifferential characterization of weakly convex functions. Recall that $\overline{\mathbb{R}}$ denotes the extended real line, i.e., $\overline{\mathbb{R}}:=\mathbb{R}\cup\{\pm\infty\}$.
We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is proper if $f(x)<\infty$ for at least one $x\in\mathbb{R}^{n}$ and $\infty<f(x)$ for all $x\in\mathbb{R}^{n}.$ We say that a function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is closed if the epigraph $\text{epi}f:\{(x,\alpha)\in\mathbb{R}^{n}\times\mathbb{R}\mid f(x)\leq\alpha\}$ is closed.
Definition 2.
For a closed function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ and a point $x$ such that $f(x)$ is finite, we define the Fréchet subdifferential at the point $x$ as
$\hat{\partial}f(x)=\mathopen{}\mathclose{{\left\{s\in\mathbb{R}^{n}\mid\liminf_{y\to x}\frac{f(y)-f(x)-\mathopen{}\mathclose{{\left\langle s,y-x}}\right\rangle}{\|y-x\|}\geq 0}}\right\}.$
Definition 3.
A function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ is called $\rho$-weakly convex if $f+\frac{\rho}{2}\|\cdot\|^{2}$ is convex.
For this class of $\rho$-weakly convex functions, the Fréchet subdifferential always exists for any point $x\in\mathbb{R}^{n}$. This is because the Fréchet subdifferential is the same as the Clarke subdifferential for $\rho$-weakly convex functions and the Clarke subdifferential always exists [li2020understanding, Section 3 and fact 5]. The class of $\rho$-weakly convex functions also has the following nice characterizations. These characterizations offer favorable properties for $\rho$-weakly convex functions, which can be viewed as extensions from strongly convex functions.
Lemma A.1 (Subdifferential characterization [davis2019stochastic, Lemma 2.1]).
The following statements are equivalent for any closed and proper function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$.
  • The function $f$ is $\rho$-weakly convex.
  • The approximate secant inequality holds:
    $f(\lambda x+(1-\lambda)y)\leq\lambda f(x)+(1-\lambda)f(y)+$ $\frac{\rho\lambda(1-\lambda)}{2}\|x-y\|^{2},\;\forall x,y\in\mathbb{R}^{n},\lambda\in[0,1].$
  • The subgradient inequality holds:
    $f(y)\geq f(x)+\mathopen{}\mathclose{{\left\langle v,y-x}}\right\rangle-\frac{\rho}{2}\|y-x\|^{2},\quad\forall x,y\in\mathbb{R}^{n},v\in\hat{\partial}f(x).$ (29)
  • The subdifferential map is hypomonotone:
    $\mathopen{}\mathclose{{\left\langle v-w,x-y}}\right\rangle\geq-\rho\|x-y\|^{2},\quad\forall x,y\in\mathbb{R}^{n},v\in\hat{\partial}f(x),\penalty 10000\ \text{and}\penalty 10000\ w\in\hat{\partial}f(y).$
If $f$ is $\mathcal{C}^{2}$-smooth, then the four properties above are all equivalent to
$\nabla^{2}f(x)\succeq-\rho I,\forall x\in\mathbb{R}^{n}.$
The subgradient inequality 29 is particularly useful in our proof for Theorem 3.1. The proof repeatedly applies 29 to establish the connection between subdifferential and cost value gap.
Appendix B Proofs in Section 2
B.1 Proof of Theorem 2.1
Recall that we say a differentialable function $f:\mathbb{R}^{n}\to\mathbb{R}$ is $L$-smooth if it satisfies
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|,\;\;\forall x,y\in\mathbb{R}^{n}.$ (30)
A differentialable $L$-smooth function $f$ also satisfy the following useful inequalities
$f(y)\leq f(x)+\mathopen{}\mathclose{{\left\langle\nabla f(x),y-x}}\right\rangle+\frac{L}{2}\|y-x\|^{2},\;\;\forall x,y\in\mathbb{R}^{n}.$ (31)
Assume that RSI 2 holds with constant $\mu_{\mathrm{r}}>0$. We first note that $\mu_{\mathrm{r}}\leq L$. Indeed, this can be seen from the following fact: let $x\in\mathbb{R}^{n},x\notin S,\hat{x}\in\Pi_{S}(x)$, we have
$\mu_{\mathrm{r}}\cdot\mathrm{dist}^{2}(x,S)=\mu_{\mathrm{r}}\|x-\hat{x}\|^{2}$ $\leq\mathopen{}\mathclose{{\left\langle\nabla f(x),x-\hat{x}}}\right\rangle=\mathopen{}\mathclose{{\left\langle\nabla f(x)-\nabla f(\hat{x}),x-\hat{x}}}\right\rangle$
$\leq L\|x-\hat{x}\|^{2}=L\cdot\mathrm{dist}^{2}(x,S),$
where the first inequality uses RSI 2, and the last inequality applies Cauchy-Schwarz and 30.
Recall that for a closed set $S\subseteq\mathbb{R}^{n}$, we denote the distance of a point $x\in\mathbb{R}^{n}$ to $S$ as $\mathrm{dist}(x,S):=\min_{y\in S}\|x-y\|$ and the projection of $x$ onto $S$ as $\Pi_{S}(x)=\operatorname*{\arg\!\min}_{y\in S}\|x-y\|$. In the following, we use $S$ to denote the set of minimizers of $f$ and let $\hat{x}_{k}$ be an element in $\Pi_{S}(x_{k})$. We first have the following inequality in terms of the distance to the solution set
$\mathrm{dist}^{2}(x_{k+1},S)$ $\leq\|x_{k}-t_{k}\nabla f(x_{k})-\hat{x}_{k}\|^{2}$ (32a)
$=\|x_{k}-\hat{x}_{k}\|^{2}-2t_{k}\mathopen{}\mathclose{{\left\langle\nabla f(x_{k}),x_{k}-\hat{x}_{k}}}\right\rangle+t_{k}^{2}\|\nabla f(x_{k})\|^{2}$ (32b)
$\leq\mathrm{dist}^{2}(x_{k},S)-2t_{k}\mu_{\mathrm{r}}\mathrm{dist}^{2}(x_{k},S)+t_{k}^{2}L^{2}\mathrm{dist}^{2}(x_{k},S)$ (32c)
$=\mathopen{}\mathclose{{\left(1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}}}\right)\mathrm{dist}^{2}(x_{k},S),$ (32d)
where 32a plugs in definition of $x_{k+1}$ and uses the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\hat{x}_{k}\|$, 32b simply expends the square, and 32c uses 2 to upper bound the inner product term and applies the $L$-smoothness inequality 30 to $\|\nabla f(x_{k})\|=\|\nabla f(x_{k})-\nabla f(\hat{x}_{k})\|$ $\leq L\|x_{k}-\hat{x}_{k}\|$ (noting that $\nabla f(\hat{x}_{k})=0$).
On the other hand, suppose the PL property 3 holds with constant $\mu_{\mathrm{q}}>0$. First, we have
$f(x_{k+1})-f(x_{k})$ $\leq\mathopen{}\mathclose{{\left\langle\nabla f(x_{k}),x_{k+1}-x_{k}}}\right\rangle+\frac{L}{2}\|x_{k+1}-x_{k}\|^{2}$
$=-t_{k}\mathopen{}\mathclose{{\left\langle\nabla f(x_{k}),\nabla f(x_{k})}}\right\rangle+\frac{Lt_{k}^{2}}{2}\|\nabla f(x_{k})\|^{2}$
$=\mathopen{}\mathclose{{\left(\frac{-2t_{k}+Lt_{k}^{2}}{2}}}\right)\|\nabla f(x_{k})\|^{2}.$
where the first inequality applies 31. If the step size $t_{k}$ is chosen such that $0<t_{k}<\frac{2}{L}$, then $\frac{-2t_{k}+Lt_{k}^{2}}{2}<0$. Then, applying PL 3 to the above inequality and subtracting $f^{\star}$ from both sides, we arrive at the linear convergence in the cost value gap
$f(x_{k+1})-f^{\star}\leq(1+\mathopen{}\mathclose{{\left(-2t_{k}+Lt_{k}^{2}}}\right)\mu_{\mathrm{p}})(f(x_{k})-f^{\star}).$ (33)
To summarize, if we choose the step size $t_{k}$ such that the constants in 32d and 33 less than one (equivalently, $0\leq 1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}<1$ and $0<t_{k}<\frac{2}{L}$), linear convergence for both the distance and the cost value gap is guaranteed. In particular, choosing $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$, we have $1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=1-\frac{\mu_{\mathrm{r}}^{2}}{L^{2}}<1$ and $t_{k}\leq\frac{L}{L^{2}}<\frac{2}{L}$ as $\mu_{\mathrm{r}}\leq L$, leading to the desired linear convergence:
$\mathrm{dist}^{2}(x_{k+1},S)$ $\leq(1-\mu_{\mathrm{r}}^{2}/L^{2})\mathrm{dist}^{2}(x_{k},S),$
$f(x_{k+1})-f^{\star}$ $\leq\frac{L^{3}+(-2\mu_{\mathrm{r}}L+\mu_{\mathrm{r}}^{2})\mu_{\mathrm{p}}}{L^{3}}(f(x_{k})-f^{\star}).$
Note that if 3 holds with $\mu_{\mathrm{p}}>0$, then $\mu_{\mathrm{p}}\leq\frac{L^{3}}{2\mu_{\mathrm{r}}L-\mu_{\mathrm{r}}^{2}}$. Otherwise, the above cost value decrease would contradict the fact that the cost value gap is always nonnegative. This choice of step size $t_{k}=\frac{\mu_{\mathrm{r}}}{L^{2}}$ is not new, which was used in [guille2022gradient, Proposition 1] or implicitly [zhang2020new, Proposition 1].
Remark 2.
Note that Theorem 2.1 is consistent with the linear convergence result in [karimi2016linear, Theorem 1] since the step size $t_{k}=\frac{1}{L}$ also satisfies $0<t_{k}<\frac{2}{L}$. Despite choosing $t_{k}=\frac{1}{L}$ renders a cleaner reduction constant $(1-\frac{\mu_{\mathrm{r}}}{L})$ in [karimi2016linear, Theorem 1], it is unclear if the distance also converges linearly under this stepsize choice (i.e., it is not guaranteed that $t_{k}=\frac{1}{L}$ leads to the constant $1-2t_{k}\mu_{\mathrm{r}}+t_{k}^{2}L^{2}=2-2\frac{\mu_{\mathrm{r}}}{L}<1$).
B.2 Proof of Lemma 2.1
We note that the following inequality holds for any $\bar{x}\in\mathbb{R}^{n}$ with a finite function value $f(\bar{x})$ [drusvyatskiy2021nonsmooth, Section 2]
$|\nabla f|(\bar{x})\leq\mathrm{dist}(0,\hat{\partial}f(\bar{x})).$ (34)
As the slope lacks basic lower-semicontinuity properties, it is important to define the limiting slope
$\overline{|\nabla f|}(\bar{x})=\liminf_{x\to\bar{x},f(x)\to f(\bar{x})}|\nabla f|(x).$ (35)
The result from [ioffe2017variational, Proposition 8.5] has shown that
$\overline{|\nabla f|}(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x})),$ (36)
where $\partial f(\bar{x})$ is the limiting subdifferential [li2020understanding, Equation 18] at $\bar{x}$ defined as
$\partial f(\bar{x})=\{s\in\mathbb{R}^{n}\mid\exists x_{k}\to\bar{x}\penalty 10000\ \text{and}\penalty 10000\ s_{k}\in\hat{\partial}f(x_{k})\penalty 10000\ \text{such that}\penalty 10000\ s_{k}\to s\}.$
Then we claim that for the class of functions whose Fréchet subdifferential and limiting subdifferential coincide, we have $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x})).$ Indeed, suppose $\hat{\partial}f(\bar{x})=\partial f(\bar{x})$, using 36 and 34, we have
$\overline{|\nabla f|}(\bar{x})=\mathrm{dist}(0,\partial f(\bar{x}))=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))\geq|\nabla f|(\bar{x}).$
On the other hand, $\overline{|\nabla f|}(\bar{x})\leq|\nabla f|(\bar{x})$ always holds true by defintion 35. Thus, $|\nabla f|(\bar{x})=\mathrm{dist}(0,\hat{\partial}f(\bar{x}))$.
Below, we show that Fréchet subdifferential and limiting subdifferential match for 1) closed and smooth; 2) closed and convex; 3) closed and weakly convex functions. We also assume that the function is proper. We first show that the result holds for a $\rho$-weakly convex closed function $f:\mathbb{R}\to\overline{\mathbb{R}}$, and the result for a proper convex closed function follows by setting $\rho=0$. By the definition of $\partial f(\bar{x})$, it is clear that $\hat{\partial}f(\bar{x})\subseteq\partial f(\bar{x})$ (choosing $x_{k}=\bar{x}$ and $s_{k}=s\in\hat{\partial}f(\bar{x})$). On the other hand, if $s\in\partial f(\bar{x})$, then we can find $\{x_{k}\}\to\bar{x},\{s_{k}\}\to s,$ and $s_{k}\in\hat{\partial}f(x_{k})$. Thus, for all $y\in\mathbb{R}^{n}$, we have
$\mathopen{}\mathclose{{\left\langle s,y-\bar{x}}}\right\rangle-\frac{\rho}{2}\|y-\bar{x}\|^{2}$ $=\lim_{k\rightarrow+\infty}\mathopen{}\mathclose{{\left\langle s_{k},y-x_{k}}}\right\rangle-\frac{\rho}{2}\|y-x_{k}\|^{2}$
$=\liminf_{k\rightarrow+\infty}\mathopen{}\mathclose{{\left\langle s_{k},y-x_{k}}}\right\rangle-\frac{\rho}{2}\|y-x_{k}\|^{2}$
$\overset{(a)}{\leq}\liminf_{k\rightarrow+\infty}f(y)-f(x_{k})$
$\overset{(b)}{\leq}f(y)-f(\bar{x}),$
where $(a)$ applies the subgradient inequality 29, and $(b)$ uses the closeseness of $f$ (i.e., lower semicontinuity $\liminf_{y\to\bar{x}}f(y)\geq f(\bar{x})$). Thus, Fréchet subdifferential and limiting subdifferential match for $\rho$-weakly convex closed functions.
For a smooth function $f$, it is also clear that $\{\nabla f(\bar{x})\}\subseteq\partial f(\bar{x})$ since $\{\nabla f(\bar{x})\}\subseteq\hat{\partial}f(\bar{x})\subseteq\partial f(\bar{x})$. On the other hand, let $s\in\partial f(\bar{x})$. We have $x_{k}\to\bar{x}$ and $\{\nabla f(x_{k})\}\to s.$ Since $f$ is smooth, the gradient $\nabla f$ is continuous, implying $\{\nabla f(x_{k})\}\to\nabla f(\bar{x})$. Thus, we have $s=\nabla f(x)$. This proves that $\{\nabla f(\bar{x})\}=\hat{\partial}f(\bar{x})=\partial f(\bar{x})$. Thus, the result follows.
Appendix C Convergence proofs of the PPM in Section 4.2
C.1 A compelete proof for Theorem 4.2
To finish the proof for 18b with coefficient $\theta_{k}\leq 1/{\sqrt{c_{k}^{2}/\mu_{\mathrm{e}}^{2}+1}}$. We first review a crucial but standard result about the proximal mapping.
Lemma C.1 ([leventhal2009metric, Equation 3.1]).
For any $c_{k}>0$ in 15, the following holds true
$\mathrm{dist}^{2}(x_{k},S)\geq\|x_{k+1}-\Pi_{S}(x_{k})\|^{2}+\|x_{k}-x_{k+1}\|^{2}.$ (37)
Proof.
As shown in [rockafellar2021characterizing], the proximal point operator for a convex function $f:\mathbb{R}^{n}\to\overline{\mathbb{R}}$ with $\alpha>0$ is firmly nonexpensive, i.e., $\forall x,y\in\mathbb{R}^{n},$
$\|\text{prox}_{\alpha f}(x)-\text{prox}_{\alpha f}(y)\|^{2}\leq\mathopen{}\mathclose{{\left\langle x-y,\text{prox}_{\alpha f}(x)-\text{prox}_{\alpha f}(y)}}\right\rangle.$ (38)
For notational convenience, let $P(x):=\text{prox}_{\alpha f}(x).$ We have
$\|x-y\|^{2}$ $=\|P(x)+(x-P(x))-(P(y)+(y-P(y)))\|^{2}$
$=\|P(x)-P(y)+(x-P(x))-(y-P(y))\|^{2}$
$=\|P(x)-P(y)\|^{2}+\|(x-P(x))-(y-P(y))\|^{2}$
$\quad+2\mathopen{}\mathclose{{\left\langle P(x)-P(y),(x-P(x))-(y-P(y))}}\right\rangle$
$\geq\|P(x)-P(y)\|^{2}+\|(x-P(x))-(y-P(y))\|^{2},$
where the last inequality comes from the fact that the inner product is nonnegative by 38.
Setting $x=x_{k}$ and $y=\Pi_{S}(x_{k})$, we have $x_{k+1}=P(x_{k})$, and $y=P(y)$, which leads to the desired result in 37. ∎
We are ready to finish the proof of Theorem 4.2. With Lemma C.1, it follows that
$\mathrm{dist}^{2}(x_{k+1},S)$ $\leq\|x_{k+1}-\Pi_{S}(x_{k})\|^{2}$
$\leq\|x_{k}-\Pi_{S}(x_{k})\|^{2}-\|x_{k}-x_{k+1}\|^{2}$
$\leq\mathrm{dist}^{2}(x_{k},S)-c_{k}^{2}\mathrm{dist}^{2}(0,\partial f(x_{k+1}))$
$\leq\mathrm{dist}^{2}(x_{k},S)-\frac{c_{k}^{2}}{\mu_{\mathrm{e}}^{2}}\mathrm{dist}^{2}(x_{k+1},S),$
where the second inequality is from 37, the third inequality comes from the optimality condition 19, and the last inequality uses EB. Rearranging terms, we have
$\mathrm{dist}^{2}(x_{k+1},S)\leq\frac{\mu_{\mathrm{e}}^{2}}{c_{k}^{2}+\mu_{\mathrm{e}}^{2}}\mathrm{dist}^{2}(x_{k},S).$
Taking a square root yields the desired result.
Appendix D Convergence proofs for the inexact PPM in Section 5
D.1 Proofs of Theorem 5.2
The optimality condition of 23 and criteria $\text{A}^{\prime}$ implies that there exist $v_{k}\in\partial f(x_{k+1})$ and $\theta_{k}\in\mathbb{R}^{n}$ such that
$0=v_{k}+(x_{k+1}-x_{k}+\theta_{k})/c_{k},\quad\|\theta_{k}\|\leq\epsilon_{k}.$
By the definition of subdifferential for convex functions, we know
$f(x^{\star})$ $\geq f(x_{k+1})+\mathopen{}\mathclose{{\left\langle-\frac{1}{c_{k}}(x_{k+1}-x_{k}+\theta_{k}),x^{\star}-x_{k+1}}}\right\rangle$
$\geq f(x_{k+1})+\frac{1}{c_{k}}\mathopen{}\mathclose{{\left\langle x_{k}-x_{k+1},x^{\star}-x_{k+1}}}\right\rangle-\frac{1}{c_{k}}\|\epsilon_{k}\|\|x^{\star}-x_{k+1}\|.$ (39)
On the other hand, we have
$\|x_{k+1}-x^{\star}\|^{2}$ $=\|x_{k}-x^{\star}\|^{2}-\|x_{k}-x_{k+1}\|^{2}+2\mathopen{}\mathclose{{\left\langle x_{k+1}-x_{k},x_{k+1}-x^{\star}}}\right\rangle$
$\leq\|x_{k}-x^{\star}\|^{2}+2\mathopen{}\mathclose{{\left\langle x_{k+1}-x_{k},x_{k+1}-x^{\star}}}\right\rangle,$ (40)
where the first identity follows from simple algebraic manipulations $\|x_{k}-x^{\star}\|^{2}=\|x_{k}-x_{k+1}+x_{k+1}-x^{\star}\|^{2}=\|x_{k}-x_{k+1}\|^{2}-2\langle x_{k+1}-x_{k},x_{k+1}-x^{\star}\rangle+\|x_{k+1}-x^{\star}\|^{2}$, and the second inequality drops a nonnegative term. Combining (39) with (40) leads to
$\|x_{k+1}-x^{\star}\|^{2}\leq\|x_{k}-x^{\star}\|^{2}-2c_{k}(f(x_{k+1})-f(x^{\star}))+2\epsilon_{k}\|x^{\star}-x_{k+1}\|.$
This improvement holds for all $k\geq 0$. Summing above inequality from 0 to k1k-1 yields
$\|x_{k}-x^{\star}\|^{2}\leq\|x_{0}-x^{\star}\|^{2}-2\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f(x^{\star}))+\sum_{j=0}^{k-1}2\epsilon_{j}\|x^{\star}-x_{j+1}\|$
This immediately leads to
$2\sum_{j=0}^{k-1}c_{j}\min_{i=0,\ldots,k}f(x_{i})-f(x^{\star})$ $\leq 2\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f(x^{\star}))\leq\|x_{0}-x^{\star}\|^{2}+2D\sum_{j=0}^{k-1}\epsilon_{j}.$
Dividing both sides by $2\sum_{j=0}^{k-1}c_{j}$ and letting $x^{\star}=\Pi_{S}(x_{0})$ gets the desired inequality. The convergence of the iterate $\{x_{k}\}$ and the boundedness of the diameter $D$ is guaranteed by Theorem 5.1.
Remark 3 (Average/weighted iterate).
Theorem 5.2 provides a convergence guarantee for the best iterate as opposed to that for the last iterate in Theorem 4.1. The guarantee for the average $\bar{x}_{k}=\frac{1}{k}\sum_{j=1}^{k}x_{j}$ or the weighted average $\tilde{x}_{k}=\frac{\sum_{j=0}^{k-1}c_{j}x_{j+1}}{\sum_{j=0}^{k-1}c_{j}}$ can also be shown using the inequality
$f(\bar{x}_{k})-f^{\star}\penalty 10000\ \text{and}\penalty 10000\ f(\tilde{x}_{k})-f^{\star}\leq\frac{\sum_{j=0}^{k-1}c_{j}(f(x_{j+1})-f^{\star})}{\sum_{j=0}^{k-1}c_{j}}.$
The proof in Theorem 5.2 borrows some proof idea from [lu2023unified, Theorem 3] with the fact that $\text{A}^{\prime}$ implies the existence of $\theta_{k}\in\mathbb{R}^{n}$ such that $\mathrm{dist}(0,H_{k}(x_{k+1}))=\frac{\|\theta_{k}\|}{c_{k}}\leq\epsilon_{k}$. ∎
D.2 Proof of Lemma 5.1
The proof is an adaptation of [luque1984asymptotic, Section 2]. For completeness, we reprove it here. Fix a $x\in\mathbb{R}^{n}$. From the update 15, we know
$f(\mathrm{prox}_{c_{k}f}(x))+\frac{1}{2c_{k}}\|\mathrm{prox}_{c_{k}f}(x)-x\|^{2}\leq f(\Pi_{S}(x))+\frac{1}{2c_{k}}\|\Pi_{S}(x)-x\|^{2}$
[…]
where the last implication is due to $f^{\star}-f(\mathrm{prox}_{c_{k}f}(x))\leq 0$. On the other hand, using triangle inequality, the nonexpensiveness of projection onto a convex set, and the above inequality, we have
$\forall x\in\mathbb{R}^{n},\penalty 10000\ \|x-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ $\leq\|x-\Pi_{S}(x)\|+\|\Pi_{S}(x)-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq\mathrm{dist}(x,S)+\|x-\text{prox}_{c_{k}f}(x)\|$
$\leq 2\mathrm{dist}(x,S).$ (41)
Also, we have
$\penalty 10000\ \|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\penalty 10000\ \|x_{k+1}-\text{prox}_{c_{k}f}(x_{k})\|+\|\text{prox}_{c_{k}f}(x_{k})-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$\leq$ $\penalty 10000\ \delta_{k}\|x_{k+1}-x_{k}\|+\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S)$
$\leq$ $\penalty 10000\ \delta_{k}\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|+\delta_{k}\|x_{k}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
$+\mathrm{dist}(\text{prox}_{c_{k}f}(x_{k}),S),$
where the second inequality is from the stopping criterion B. As $\delta_{k}\to 0$, choosing $\hat{k}$ such that $\delta_{k}<1$ for all $k\geq\hat{k}$, we have
$(1-\delta_{k})\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$
[…]
Finally, using the fact $\mathrm{dist}(x_{k+1},S)\leq\|x_{k+1}-\Pi_{S}(\text{prox}_{c_{k}f}(x))\|$ and 41 for $x=x_{k}$ completes the proof.
Declarations
[…]